Reports 1-1 of 1 Clear search Modify search
MIF (ITF Control)
yuta.michimura - 11:48 Tuesday 13 August 2024 (30825) Print this report
f1 and 1st higher order mode are very close in the arm cavities

[Haoyu Wang, Yuta Michimura]

We found that f1 sideband and 1st higher order mode are very close and overlapping in the arm cavities.
There might be some effects in ASC.
(I'm not sure if this is known or not...)

Background:
 - As a preparation for the arm cavity round-trip loss measurement, we did the arm cavity scan to check the mode-matching (klog #30823).
 - We found that the sideband locations were consistent with Hirose+ measurement in Nov 2023 (klog #27767), but not with Enomoto+ measurements in Dec 2018 (klog #7332 and klog #7307)

What we did:
 - After aligning Xarm, we intentionally misaligned ETMX in YAW (changed K1:VIS-ETMX_TM_SET_Y_OFFSET=-18 to -22) so that we can see higher order modes better.
 - Scanned Xarm following the same steps we took last night (klog #30823).

Results:
 - Attachement #1 shows the result of the Xarm cavity scan. 1st and 2nd higher order modes can be clearly seen from the error signals and transmission peaks. f1 sideband and 1st higher order mode are overlapped (see Attachment #2 for the zoomed in view).
 - In the original design, f1 and f2 sidebands are located in between 0th and 1st order modes (Attachment #3), but if we use the measured values, HOM1 and f1 are overlapped (Attachment #4), explaining the current situation. Table below are the list of parameters used to calculate these mode spacing plots.

  Design (see OptParam wiki page) Measured (individually)
Arm length 3000 m 2999.990(2) m (klog #7332; could be old)
ITM RoC 1900 m 1904.54 +/- 2.08 m (PhysRevApplied.14.014021)
ETM RoC 1900 m 1907.83 +/- 2.14 m (PhysRevApplied.14.014021)
f1 sideband frequency 16.880962 MHz 5.624365627 MHz *3 (klog #30092)
f2 sideband frequency 45.015898 MHz 5.624365627 MHz *8 (klog #30092)


Note:
  - We confirmed that the mode-matching is pretty good when aligned.
  - Compared with the measurement in Nov 2023 (klog #27767), K1:LSC-TR_IRX_NORM_OUT_DQ seems noisy, and this is preventing us from seeing sideband beaks in the transmission. Not sure why it is noisier now (Different PD gains and/or whitening gains? Notch filters?).

Next:
 - We might need to tune IMC length and PRC length to shift the sideband frequencies.

Images attached to this report
Comments to this report:
yuta.michimura - 14:13 Wednesday 14 August 2024 (30828) Print this report

[Chiaki Hirose, Haoyu Wang, Yuta Michimura]
 

  - Compared with the measurement in Nov 2023 (klog #27767), K1:LSC-TR_IRX_NORM_OUT_DQ seems noisy, and this is preventing us from seeing sideband beaks in the transmission. Not sure why it is noisier now (Different PD gains and/or whitening gains? Notch filters?).
 


We checked the past data and confirmed that the IMC output power was 9.8 W when Hirose-san did the cavity scan in Nov 2023 (Attachment #1), and it is now 1.2 W (Attachment #2). We also checked that the normalization factors in K1:TMS-X_IR_PDA1 and K1:LSC-TR_IRY_NORM filter banks were the same, and K1:LSC-TR_IRX_NORM_OUT_DQ is roughly one when Xarm was locked with PRM mis-aligned, suggesting that PD gains and whitening filter gains were the same (0dB for PD setting and no whitening gain). These explain why the noise level was about 10 times lower in Nov 2023.

 

Images attached to this comment
takafumi.ushiba - 9:36 Thursday 15 August 2024 (30832) Print this report

[Michimura, Ushiba]

Some discussions related to this issue.

Current IMC FSR (5.6243656 MHz, klog #30092) is far from the nomina value (5.6269873 MHz, MIF params) designed by MIF.
Somehow, IMC length written in IOO params (53.300 m) is several cm different from the values calculated from designed FSR (53.278 m).
Sideband frequency of IMC is now somehow 13.77 MHz (klog30602) while it was 13.78 MHz (klog23038). We are not so sure when this change happened.

Search Help
×

Warning

×